A manufacturing company prepays its insurance coverage for a three-year period. The premium for the three years is $3,120 and is paid at the beginning of the first year. Ninety percent of the premium applies to manufacturing operations and ten percent applies to selling and administrative activities. What amounts should be considered product and period costs respectively for the first year of coverage?

Product Period
A) $3,120 $0
B) $2,808 $312
C) $1,872 $208
D) $936 $104
Answer #1

Calculate amount considered to product and period Cost for the first year

First year insurance = 3120/3 = 1040

Product Cost (insurance) = 1040*90% = $936

Period Cost (insurance) = 1040*10% = $104

So answer is d)

Product Cost Period Cost
D) $936 $104